You want to buy a $265,000 home. You plan to pay 20% as a down payment, and take out a 30 year loan for the rest.
a) How much is the loan amount going to be?
b) What will your monthly payments be if the interest rate is 6%? (Round up to the cent.)

c) What will your monthly payments be if the interest rate is 7%? (Round up to the cent.)

Answer by math_tutor2020(3817)


Part a)

You pay 20% as a down payment
You'll be loaned the remaining 80%
80% of 265,000 = 0.80*265,000 = 212,000

Answer: $212,000

============================================

Part b)

P = monthly payment
L = loan amount = $212,000
i = interest rate per month = 0.06/12 = 0.005
n = number of months = 12*30 = 360

P = (L*i)/( 1-(1+i)^(-n) )
P = (212,000*0.005)/( 1-(1+0.005)^(-360) )
P = 1,271.04711332384
P = 1,271.05

Answer: $1,271.05

============================================

Part c)

P = monthly payment
L = loan amount = $212,000
i = interest rate per month = 0.07/12 = 0.0058333333 (approximate)
n = number of months = 12*30 = 360

P = (L*i)/( 1-(1+i)^(-n) )
P = (212,000*0.0058333333)/( 1-(1+0.0058333333)^(-360) )
P = 1,410.44128408472
P = 1,410.45

Answer: $1,410.45